OpenStudy (anonymous):

-2x+9y=5 Find the value of y when x=11

OpenStudy (anonymous):

-2(11) +9y = 5 9y = 5 + 2(11)

OpenStudy (anonymous):

Wut?

OpenStudy (anonymous):

-2x + 9y = 5 what will be the value of y if x =11 we put x =11 in equation -2(11) +9y = 5 -22 +9y = 5 adding 22 both sides -22 +22 + 9y = 5+22 9 y = 27 y = 27/9

OpenStudy (anonymous):

solve using that equation and you find the answer is 3

OpenStudy (anonymous):

Thanks!

OpenStudy (catlover5925):